OpenStudy (anonymous):

FREE MEDAL!!! if you answer!!!! With a tailwind, a jet flew 2000 miles in 4 hours, but the return trip against the same wind required 5 hours. find the jet's speed and the wind speed.

OpenStudy (pokemon23):

Hero and moneybird got this in the bag

OpenStudy (anonymous):

Let the jet speed be a and wind speed be b (a +b) 4 = 2000 5(a-b) = 2000

hero (hero):

I don't know why everyone keeps saying "Free Medal". The medals are free anyway. No one has to pay for them.

OpenStudy (pokemon23):

I just gave you a free medals

OpenStudy (anonymous):

a + b = 500 a - b = 400 2a = 900 a = 450 b = 50

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

answererrrrrrred